OpenStudy (anonymous):

A triangle ABC is inscribed in a circle and the bisectors of the angles meet the circumference at X,Y and Z. The angles of triangle XYZ are respectively_________

OpenStudy (anonymous):

A) 90, 60, 30 B) A/2 , B/2 , C/2 C) 90-A/2 , 90-B/2 , 90-C/2 D) B/2,A/2 , A/2-B/2

OpenStudy (asnaseer):

Use the fact that an angle subtended by an arc at the circumference is always the same. i.e.:  angle PRQ = angle PSQ so:  angle AXY = angle ABY = B/2 angle AXZ = angle ACZ = C/2 therefore angle YXZ = B/2 + C/2 = (B+C)/2 follow the same procedure for the other angles. the answer I get, however, does not seem to match any of the ones you have listed, so maybe I have got the diagram wrong - can you please clarify?